# Physics-Informed Weakly Supervised Learning for Interatomic Potentials

### v1

@@ Abstract
Machine learning plays an increasingly important role in computational chemistry and materials science, complementing computationally intensive ab initio and first-principles methods. Despite their utility, machine-learning models often lack generalization capability and robustness during atomistic simulations, yielding unphysical energy and force predictions that hinder their real-world applications. We address this challenge by introducing a physics-informed, weakly supervised approach for training machine-learned interatomic potentials (MLIPs). We introduce two novel loss functions, extrapolating the potential energy via a Taylor expansion and using the concept of conservative forces. Our approach improves the accuracy of MLIPs applied to training tasks with sparse training data sets and reduces the need for pre-training computationally demanding models with large data sets. Particularly, we perform extensive experiments demonstrating reduced energy and force errors—often lower by a factor of two—for various baseline models and benchmark data sets. Finally, we show that our approach facilitates MLIPs’ training in a setting where the computation of forces is infeasible at the reference level, such as those employing complete-basis-set extrapolation.
@@ 1 Introduction
Ab initio and first-principles methods are inevitable for the computer-aided exploration of molecular and material properties used in the chemical sciences and engineering [1, 2, 3]. However, commonly employed ab initio and first-principles approaches—such as coupled cluster (CC) [4, 5] and density functional theory (DFT) [6, 7], respectively—require substantial compute resources. Thus, they typically allow only for atomistic simulations of small- to medium-sized atomic systems and restrict the accessible simulation times, which affects the accuracy of estimated molecular and material properties. Classical force fields can extend these length and time scales, providing a computationally efficient alternative to first-principles approaches, but often lack accuracy. Machine-learning-based models hold promise to bridge the gap between first-principles and classical approaches, yielding computationally efficient and accurate machine-learned interatomic potentials (MLIPs) [8, 9, 10, 11, 12, 13]. These MLIPs, however, face several challenges. They require the generation of training data sets that sufficiently cover configurational (atom positions) and compositional (atom types) spaces using, e.g., molecular dynamics simulations based on ab initio or first-principles approaches. Given the high computational cost of the commonly used data generation approaches, the resulting training data sets are often sparse and prevent the application of MLIPs to new molecular and material systems. Active learning can be used to address this challenge [14, 15, 16], but may require non-negligible computer resources. Furthermore, MLIPs often lack sufficient generalization capability and robustness during atomistic simulations, i.e., they are sensitive to outliers and local perturbations of atomic structures. This sensitivity of ML-based models is caused by existing data sets and data generation techniques not providing sufficient coverage of configurational and compositional spaces.
@@
Figure 1: Schematic illustration of physics-informed weakly supervised losses used in this work. (a) Taylor-expansion-based weak label (WL) loss with approximate labels obtained from reference energies and atomic forces [17]. (b) Physics-inspired Taylor-expansion-consistency (PITC) loss with approximate labels obtained from energies and atomic forces predicted by an MLIP. (c) Physics-inspired spatial consistency (PISC) loss with approximate labels obtained from energies and atomic forces predicted by an MLIP. Here, $E(\mathcal{S};\bm{\theta})$ and $\mathbf{F}_{i}(\mathcal{S};\bm{\theta})$ denote the potential energy and atomic forces predicted by an MLIP parametrized by $\bm{\theta}$, $\mathcal{S}$ and $\mathcal{S}_{\delta\mathbf{r}}$ define the original atomic structure and the one perturbed by $\delta\mathbf{r}$.
@@
Contributions. This paper addresses these challenges using a physics-informed weakly supervised learning (PIWSL) approach. Our method is designed to learn an MLIP, which can accurately predict the potential energy and atomistic forces for an atomic system exposed to local perturbations. In particular, our contributions are as follows: (i) We introduce PIWSL based on basic physical principles, such as the concept of conservative forces. We combine it with extrapolating the potential energy via a Taylor expansion and derive two novel physics-informed loss functions, schematically illustrated in Fig. 1 (b) and (c). Particularly, we obtain physics-informed Taylor-expansion-based consistency (PITC) and physics-informed spatial consistency (PISC) losses, which build the basis for the PIWSL approach. (ii) By conducting extensive experiments, we demonstrate that PIWSL facilitates the training of MLIPs without access to large training data sets. (iii) We also observe that PIWSL improves accuracy in predicted total energies and atomic forces, even without access to force labels. This scenario is expected when training MLIPs with reference methods for which calculating atomic forces is infeasible [18, 19, 20]. Thus, our results open new possibilities for training MLIPs using highly accurate energy labels, such as those obtained by extrapolating CCSD(T) energies to the complete basis set (CBS) limit [21, 22]. (iv) Finally, PIWSL mitigates sensitivity issues associated with limited sizes of available data sets by taking into account the potential energy response to local perturbations in atomic structures.
@@ 2 Related Work
Machine-Learned Interatomic Potentials. There is a growing interest in using ML-based models for investigating molecular and material systems as they allow performing atomistic simulations with an accuracy on par with first-principles methods but at a fraction of the computational cost. The field of machine-learned interatomic potentials (MLIPs) emerged over two decades ago [23] and has been one of the most active research directions since then [24, 25, 26, 8, 27, 28, 29, 30, 31, 32, 33, 29, 34, 35, 36, 37, 38, 39, 40]. The development of local higher-body-order representations [27, 31, 32, 33] and the emergence of equivariant message-passing neural networks (MPNNs) [29, 34, 35, 36, 37, 38, 39, 40] significantly advanced the field. These methods enable the cost-efficient generation of accurate MLIPs for modeling interactions in many-body atomic systems and account for crucial inductive biases as the invariance of the potential energy under rotation.
Physics-Informed Machine Learning. Physics-informed ML aims to model physical systems using data-driven techniques and incorporates physics principles into ML-based models. For example, MLIPs based on equivariant MPNNs enforce the invariance of the potential energy under rotation and use equivariant features to enrich the building of many-body contributions to it [29, 38, 40, 39, 37]. Furthermore, physics constraints can be integrated via auxiliary loss functions, prompting ML models to learn important physical relationships, as demonstrated for physics-informed neural networks (PINNs) [41, 42], which learn to model solutions of partial differential equations by minimizing residuals during training. Applying physics-informed ML to molecular modeling has gained attraction in both ML and computational chemistry communities [43, 44]. As such, prior work [17] has motivated our current research and is discussed in more detail in subsequent sections.
@@ 3 Background and Problem Definition
Machine-Learned Interatomic Potentials. An atomic configuration, denoted as $\mathcal{S}=\{\mathbf{r}_{i},Z_{i}\}_{i=1}^{N_{\mathrm{at}}}$, contains $N_{\mathrm{at}}$ atoms and is defined by atom positions $\mathbf{r}_{i}\in\mathbb{R}^{3}$ and atom types $Z_{i}\in\mathbb{N}$. We consider mapping atomic configurations to scalar energies, i.e., $f_{\bm{\theta}}:\mathcal{S}\mapsto E\in\mathbb{R}$ with $\bm{\theta}$ denoting trainable parameters. We define $E(\mathcal{S};\bm{\theta})$ as the energy predicted by an MLIP for an atomic configuration $\mathcal{S}$. For most MLIPs, atomic forces are computed as the negative gradients of the potential energy with respect to atom positions, i.e., $\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)=-\nabla_{\mathbf{r}_{i}}E\left(\mathcal{S};\bm{\theta}\right)$. In this way, these MLIPs ensure that the resulting forces are conservative (curl-free) and the total energy is conserved during a dynamic simulation. However, some models are designed to predict atomic forces directly [45, 36, 37, 9]. While this approach avoids expensive gradient computations, it violates the law of energy conservation [46].
@@
$\mathcal{L}\left(\mathcal{D};\bm{\theta}\right)=\sum_{\mathcal{S}\in\mathcal{D}}L\left(S;\bm{\theta}\right)=\sum_{\mathcal{S}\in\mathcal{D}}\Big{[}C_{\mathrm{e}}\ell\left(E\left(\mathcal{S};\bm{\theta}\right),E^{\mathrm{ref}}_{\mathcal{S}}\right)+C_{\mathrm{f}}\sum_{i=1}^{N_{\mathrm{at}}}\ell\left(\mathbf{F}_{i}(\mathcal{S};\bm{\theta}),\mathbf{F}_{i,\mathcal{S}}^{\mathrm{ref}}\right)\Big{]}.$ (1)
@@
Weakly Supervised Learning. Generating many reference labels with a first-principles approach is challenging due to the high computational cost. Furthermore, the calculation of atomic forces can be infeasible for some high-accuracy ab initio methods, e.g., for CCSD(T)/CBS. In this work, we focus on weakly supervised learning methods to improve the performance of MLIPs in scenarios when only a limited amount of data is available. These involve the generation of approximate but physically motivated total energies for atomic structures generated by small perturbations of their atomic positions, i.e., $\mathcal{S}_{\delta\mathbf{r}}=\{\mathbf{r}_{i}+\delta\mathbf{r}_{i},Z_{i}\}_{i=1}^{N_{\mathrm{at}}}$ with a perturbation vector $\delta\mathbf{r}$ where $\delta\mathbf{r}_{i}$ is the perturbation vector for atom $i$. Approximate labels are computed with MLIPs during their training.
@@ 4 Physics-informed Weakly Supervised Learning
For MLIPs, the generation of approximate labels employed in weakly supervised losses is highly non-trivial. Small perturbationss in atomic structures can lead to significant changes in energies and atomic forces. Thus, standard approaches that are effective for many ML tasks [47] are typically not applicable to MLIPs. To address this problem, we propose a physics-informed weakly supervised learning approach that involves (i) a Taylor expansion of the potential energy for computing the response to atomic perturbations and (ii) spatial consistency to estimate the displaced potential energy based on the concept of conservative forces. We finally introduce the PIWSL loss term, combining both classes of weakly supervised loss functions with the supervised loss.
@@ 4.1 Physics-Informed Taylor-Expansion-Based Consistency Loss
This section introduces the physics-informed Taylor-expansion-based consistency (PITC) loss. Particularly, we relate the energy predicted directly for a displaced atomic configuration with the energy obtained by the Taylor expansion from the original configuration; see Fig. 1 (b). We estimate the energy for an atomic structure $\mathcal{S}$ drawn from the training data set with atomic positions displaced by a vector $\delta\mathbf{r}$: $\mathcal{S}_{\delta\mathbf{r}}=\{\mathbf{r}_{i}+\delta\mathbf{r}_{i},Z_{i}\}_{i=1}^{N_{\mathrm{at}}}$. For this atomic configuration, we expand the energy predicted by an MLIP in its first-order Taylor series around the atomic perturbation vector $\delta\mathbf{r}_{i}$ and obtain
$\begin{split}E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)&\approx E\left(\mathcal{S};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)\rangle+\mathcal{O}\left(\|\delta\mathbf{r}\|^{2}\right),\end{split}$ (2)
@@
$\begin{split}L_{\mathrm{PITC}}\left(\mathcal{S};\bm{\theta}\right)&=\ell\Big{(}E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right),E\left(\mathcal{S};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)\rangle\Big{)},\end{split}$ (3)
where $\ell$ denotes a point-wise loss for regression problems and $\delta\bm{r}$ is randomly-sampled or determined adversarially; see Section 4.4 for more details. Hence, whenever we encounter a structure $\mathcal{S}$ in a batch during training, a new $\delta\bm{r}$ is computed for each $\mathcal{S}$.

### v2

@@ Abstract
Machine learning plays an increasingly important role in computational chemistry and materials science, complementing computationally intensive ab initio and first-principles methods. Despite their utility, machine-learning models often lack generalization capability and robustness during atomistic simulations, yielding unphysical energy and force predictions that hinder their real-world applications. We address this challenge by introducing a physics-informed, weakly supervised approach for training machine-learned interatomic potentials (MLIPs). We introduce two novel loss functions, extrapolating the potential energy via a Taylor expansion and using the concept of conservative forces. Our approach improves the accuracy of MLIPs applied to training tasks with sparse training data sets and reduces the need for pre-training computationally demanding models with large data sets. Particularly, we perform extensive experiments demonstrating reduced energy and force errors—often lower by a factor of two—for various baseline models and benchmark data sets. Moreover, we demonstrate improved robustness during MD simulations of the MLIP models trained with the proposed weakly supervised loss. Finally, our approach improves the fine-tuning of foundation models on sparse, highly accurate ab initio data. An implementation of our method and scripts for executing experiments are available at https://github.com/nec-research/PICPS-ML4Sci.
Interatomic Potential Model, Physics Informed method, Weakly Supervised Approach, Machine Learning
@@ 1 Introduction
Ab initio and first-principles methods are inevitable for the computer-aided exploration of molecular and material properties used in the chemical sciences and engineering (Parrinello, 1997; Carloni et al., 2002; Iftimie et al., 2005). However, commonly employed ab initio and first-principles approaches—such as coupled cluster (CC) (Purvis & Bartlett, 1982; Bartlett & Musiał, 2007) and density functional theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965), respectively—require substantial compute resources. Thus, they typically allow only for atomistic simulations of small- to medium-sized atomic systems and restrict the accessible simulation times, which affects the accuracy of estimated molecular and material properties. Classical force fields can extend these length and time scales, providing a computationally efficient alternative to first-principles approaches, but often lack accuracy. Machine-learning-based models hold promise to bridge the gap between first-principles and classical approaches, yielding computationally efficient and accurate machine-learned interatomic potentials (MLIPs) (Smith et al., 2017; Chanussot et al., 2021; Unke et al., 2021; Merchant et al., 2023; Kovács et al., 2023; Batatia et al., 2023).
These MLIPs, however, face several challenges. They require the generation of training data sets that sufficiently cover configurational (atom positions) and compositional (atom types) spaces using, e.g., molecular dynamics (MD) simulations based on ab initio or first-principles approaches. Given the high computational cost of the commonly used data generation approaches, the resulting training data sets are often sparse and prevent the application of MLIPs to new molecular and material systems. Active learning can be used to address this challenge (Li et al., 2015; Vandermause et al., 2020; Zaverkin & Kästner, 2021; Zaverkin et al., 2022; van der Oord et al., 2023; Zaverkin et al., 2024b), but still requires generating a non-negligible number of first-principles (e.g., DFT) data to train the initial model, which is then used to explore the phase space with sufficiently long MD simulations. Hence, a strong motivation for the proposed method is its use in combination with active learning to acquire additional training data. Furthermore, MLIPs often lack sufficient generalization capability and robustness during MD simulations, i.e., they are sensitive to outliers and local perturbations of atomic structures. This sensitivity of ML models is caused by existing data sets and data generation techniques not providing sufficient coverage of configurational and compositional spaces.
@@
Figure 1: Schematic illustration of physics-informed weakly supervised losses used in this work. (a) Taylor-expansion-based weak label (WL) loss with approximate labels obtained from reference energies and atomic forces (Cooper et al., 2020). (b) Physics-inspired Taylor-expansion-consistency (PITC) loss with approximate labels obtained from energies and atomic forces predicted by an MLIP. (c) Physics-inspired spatial consistency (PISC) loss with approximate labels obtained from energies and atomic forces predicted by an MLIP. Here, $E(\mathcal{S};\bm{\theta})$ and $\mathbf{F}_{i}(\mathcal{S};\bm{\theta})$ denote the potential energy and atomic forces predicted by an MLIP parametrized by $\bm{\theta}$, $\mathcal{S}$ and $\mathcal{S}_{\delta\mathbf{r}}$ define the original atomic structure and the one perturbed by $\delta\mathbf{r}$.
@@
Contributions. This paper addresses these challenges using a physics-informed weakly supervised learning (PIWSL) approach. Our method is designed to learn an MLIP, which can accurately predict the potential energy and atomistic forces for an atomic system exposed to local perturbations. In particular, our contributions are as follows: (i) We introduce PIWSL based on basic physical principles, such as the concept of conservative forces. We combine it with extrapolating the potential energy via a Taylor expansion and derive two novel physics-informed loss functions, schematically illustrated in figure 1 (b) and (c). Particularly, we obtain physics-informed Taylor-expansion-based consistency (PITC) and physics-informed spatial consistency (PISC) losses, which build the basis for the PIWSL approach. (ii) By conducting extensive experiments, we demonstrate that PIWSL facilitates the training of MLIPs without access to large training data sets. Furthermore, we show that MLIPs trained with PIWSL are more robust during MD simulations compared to those trained without it. (iii) We also observe that PIWSL improves accuracy in predicted total energies and atomic forces, even without access to force labels. This scenario is particularly relevant for fine-tuning foundation MLIP models with reference ab initio methods where atomic force calculations are computationally intractable or entirely unavailable (Smith et al., 2019, 2020; Zaverkin et al., 2023). Thus, our results open new possibilities for training MLIPs using highly accurate energy labels, such as those obtained by extrapolating CCSD(T) energies to the complete basis set (CBS) limit (Hobza & Šponer, 2002; Feller et al., 2006). (iv) Finally, PIWSL mitigates sensitivity issues during MD simulations associated with limited sizes of available data sets by taking into account the potential energy response to local perturbations in atomic structures.
@@ 2 Related Work
Machine-Learned Interatomic Potentials. There is a growing interest in using ML-based models for investigating molecular and material systems as they allow performing atomistic simulations with an accuracy on par with first-principles methods but at a fraction of the computational cost. The field of machine-learned interatomic potentials (MLIPs) emerged over two decades ago (Blank et al., 1995) and has been one of the most active research directions since then (Behler & Parrinello, 2007; Artrith et al., 2011; Artrith & Urban, 2016; Smith et al., 2017; Shapeev, 2016; Schütt et al., 2017; Thomas et al., 2018; Unke & Meuwly, 2019; Drautz, 2019; Zaverkin & Kästner, 2020; Zaverkin et al., 2021; Thomas et al., 2018; Schütt et al., 2021; Shuaibi et al., 2021a; Passaro & Zitnick, 2023; Liao et al., 2023; Batzner et al., 2022; Musaelian et al., 2023; Batatia et al., 2022; Zaverkin et al., 2024a). The development of local higher-body-order representations (Shapeev, 2016; Drautz, 2019; Zaverkin & Kästner, 2020; Zaverkin et al., 2021) and the emergence of equivariant message-passing neural networks (MPNNs) (Thomas et al., 2018; Schütt et al., 2021; Shuaibi et al., 2021a; Passaro & Zitnick, 2023; Liao et al., 2023; Batzner et al., 2022; Musaelian et al., 2023; Batatia et al., 2022; Zaverkin et al., 2024a) significantly advanced the field. These methods enable the cost-efficient generation of accurate MLIPs for modeling interactions in many-body atomic systems and account for crucial inductive biases as the invariance of the potential energy under rotation.
Physics-Informed Machine Learning. Physics-informed ML aims to model physical systems using data-driven techniques and incorporates physics principles into ML-based models. For example, MLIPs based on equivariant MPNNs enforce the invariance of the potential energy under rotation and use equivariant features to enrich the building of many-body contributions to it (Thomas et al., 2018; Batzner et al., 2022; Batatia et al., 2022; Musaelian et al., 2023; Liao et al., 2023; Zaverkin et al., 2024a). Furthermore, physics constraints can be integrated via auxiliary loss functions, prompting ML models to learn important physical relationships, as demonstrated for physics-informed neural networks (PINNs) (Raissi et al., 2019; Cai et al., 2022), which learn to model solutions of partial differential equations by minimizing residuals during training. Applying physics-informed ML to molecular modeling has gained attraction in both ML and computational chemistry communities (Godwin et al., 2022; Ni et al., 2024). As such, prior work (Cooper et al., 2020) has motivated our current research and is discussed in more detail in subsequent sections.
@@ 3 Background and Problem Definition
Machine-Learned Interatomic Potentials. An atomic configuration, denoted as $\mathcal{S}=\{\mathbf{r}_{i},Z_{i}\}_{i=1}^{N_{\mathrm{at}}}$, contains $N_{\mathrm{at}}$ atoms and is defined by atom positions $\mathbf{r}_{i}\in\mathbb{R}^{3}$ and atom types $Z_{i}\in\mathbb{N}$. We consider mapping atomic configurations to scalar energies, i.e., $f_{\bm{\theta}}:\mathcal{S}\mapsto E\in\mathbb{R}$ with $\bm{\theta}$ denoting trainable parameters. We define $E(\mathcal{S};\bm{\theta})$ as the energy predicted by an MLIP for an atomic configuration $\mathcal{S}$. For most MLIPs, atomic forces are computed as the negative gradients of the potential energy with respect to atom positions, i.e., $\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)=-\nabla_{\mathbf{r}_{i}}E\left(\mathcal{S};\bm{\theta}\right)$. In this way, these MLIPs ensure that the resulting forces are conservative (curl-free) and the total energy is conserved during a dynamic simulation. However, some models are designed to predict atomic forces directly (Hu et al., 2021; Passaro & Zitnick, 2023; Liao et al., 2023; Chanussot et al., 2021). While this approach avoids expensive gradient computations, it violates the law of energy conservation (Chmiela et al., 2017).
@@
$\mathcal{L}\left(\mathcal{D};\bm{\theta}\right)=\sum_{\mathcal{S}\in\mathcal{D}}L\left(S;\bm{\theta}\right)$
$=\sum_{\mathcal{S}\in\mathcal{D}}\Big{[}C_{\mathrm{e}}\ell\left(E\left(\mathcal{S};\bm{\theta}\right),E^{\mathrm{ref}}_{\mathcal{S}}\right)+C_{\mathrm{f}}\sum_{i=1}^{N_{\mathrm{at}}}\ell\left(\mathbf{F}_{i}(\mathcal{S};\bm{\theta}),\mathbf{F}_{i,\mathcal{S}}^{\mathrm{ref}}\right)\Big{]}.$ (1)
@@
Weakly Supervised Learning. Generating many reference labels with a first-principles approach is challenging due to the high computational cost. Furthermore, the calculation of atomic forces can be infeasible for some high-accuracy ab initio methods, e.g., for CCSD(T)/CBS. In this work, we focus on weakly supervised learning methods to improve the performance of MLIPs in scenarios when only a limited amount of data is available. These involve the generation of approximate but physically motivated total energies for atomic structures generated by small perturbations of their atomic positions, i.e., $\mathcal{S}_{\delta\mathbf{r}}=\{\mathbf{r}_{i}+\delta\mathbf{r}_{i},Z_{i}\}_{i=1}^{N_{\mathrm{at}}}$ with a perturbation vector $\delta\mathbf{r}$, where $\delta\mathbf{r}_{i}$ is the perturbation vector for atom $i$. Approximate labels are computed with MLIPs during their training.
@@ 4 Physics-informed Weakly Supervised Learning
For MLIPs, the generation of approximate labels employed in weakly supervised losses is highly non-trivial. Small perturbations in atomic structures can significantly change energies and atomic forces. Thus, standard approaches that are effective for many ML tasks (Yang et al., 2022) are typically not applicable to MLIPs. To address this problem, we propose a physics-informed weakly supervised learning approach that involves (i) a Taylor expansion of the potential energy for computing the response to atomic perturbations and (ii) spatial consistency to estimate the displaced potential energy based on the concept of conservative forces. We finally introduce the PIWSL loss term, combining both classes of weakly supervised loss functions with the supervised loss.
@@ 4.1 Physics-Informed Taylor-Expansion-Based Consistency Loss
This section introduces the physics-informed Taylor-expansion-based consistency (PITC) loss. Particularly, we relate the energy predicted directly for a displaced atomic configuration with the energy obtained by the Taylor expansion from the original configuration; see figure 1 (b). We estimate the energy for an atomic structure $\mathcal{S}$ drawn from the training data set with atomic positions displaced by a vector $\delta\mathbf{r}$: $\mathcal{S}_{\delta\mathbf{r}}=\{\mathbf{r}_{i}+\delta\mathbf{r}_{i},Z_{i}\}_{i=1}^{N_{\mathrm{at}}}$. For this atomic configuration, we expand the energy predicted by an MLIP in its second-order Taylor series around the atomic perturbation vector $\delta\mathbf{r}_{i}$ and obtain
$E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)\approx E\left(\mathcal{S};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\left[\langle\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)\rangle\right.$
$\left.+k_{\rm 2nd}\langle\delta\mathbf{r}_{i},\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)-\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)\rangle\right]+\mathcal{O}\left(\|\delta\mathbf{r}\|^{3}\right),$ (2)
@@
$L_{\mathrm{PITC}}\left(\mathcal{S};\bm{\theta}\right)=\ell\Big{(}E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right),E\left(\mathcal{S};\bm{\theta}\right)$
$-\sum_{i=1}^{N_{\mathrm{at}}}\langle\delta\mathbf{r}_{i},(1-k_{\rm 2nd})\mathbf{F}_{i}\left(\mathcal{S};\bm{\theta}\right)+k_{\rm 2nd}\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)\rangle\Big{)},$ (3)
where $\ell$ denotes a point-wise loss for regression problems and $\delta\bm{r}$ is randomly-sampled or determined adversarially; see section 4.4 for more details. Hence, whenever we encounter a structure $\mathcal{S}$ in a batch during training, a new $\delta\bm{r}$ is computed for each $\mathcal{S}$. Empirically, second-order terms become important when MLIP prediction errors reach the first-order accuracy level, meaning that first-order terms alone are no longer sufficient to improve accuracy or when the data primarily consists of relaxed structures.